- A1/3
- B2/3
- C3/4
- D5/12
Step-by-Step Solution
Total marbles = 12, green = 4, so P(green) = 4/12 = 1/3. By the complement rule P(not green) = 1 − 1/3 = 2/3.
Answer: B) 2/3
Don't just memorise the answer — build a daily practice habit.
Practice 5 Similar Questions Free →All Data & Probability practice questions